Nine times the difference of 10 and a number equals 135. Let x represent the number. use the given conditions to complete the following:
A.) write the equation
B.) find the number by solving the equation

Answer by TimothyLamb(4379)

9(10 - x) = 135
90 - 9x = 135
9x = 90 - 135
x = (90 - 135)/9
x = -5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
Send and receive free text messages online:
https://sooeet.com